# Order Matters in Hallucination: Reasoning Order as Benchmark and Reflexive Prompting for Large-Language-Models

### v3

@@ Abstract
Introduction
In the field of natural language processing (NLP), large language models (LLMs) have achieved superior performance by leveraging their vast number of parameters and extensive training data, surpassing traditional language models across various domains. Due to their exceptional generative and reasoning capabilities, LLMs have been widely adopted in both academia and industry, including but not limited to education (Kasneci et al. 2023), healthcare (Sallam 2023), and finance (Wu et al. 2023). The evolution of LLMs, marked by significant milestones such as the development of models like GPT-3 (Brown et al. 2020) and GPT-4 (Achiam et al. 2023), has revolutionized the way we approach language understanding and generation tasks.
However, as artificial intelligence becomes more widely adopted, its potential risks are increasingly exposed to public scrutiny. In areas with high ethical risks, the use of machine learning methods is particularly concerning due to the black-box nature of these models, which results in an opaque decision-making process. This lack of transparency makes it challenging to prevent, control, and hold accountable for the associated risks. For example, research by (Morley et al. 2020) noted that algorithmic outcomes are usually in the form of probabilities, which means infallibility cannot be assured, making it difficult to establish causal relationships.
In the realm of LLMs, although these models can provide a reasoning process for their outputs, issues such as the hallucination phenomenon (Yao et al. 2023) still persist, hindering their broader application. The hallucination problem refers to instances where models produce outputs that are grammatically and logically coherent but factually inaccurate or entirely fabricated. Recently, researchers discovered that almost all LLMs incorrectly infer 9.11$>$9.9 during numerical comparison, which has sparked widespread discussion within the community. When asked, ”Which is greater: 9.11 or 9.9?” nearly all large language models respond that 9.11 is greater, as shown in Figure 1. Furthermore, some models provide incorrect reasoning processes to justify this erroneous answer, while others present correct reasoning processes but still arrive at the wrong conclusion.
@@ Introduction
Although recent studies have demonstrated the impact of prompting strategies on mitigating hallucination problems across various fields (Si et al. 2022; Mündler et al. 2023; Ji et al. 2023; Dhuliawala et al. 2023), there is a lack of research and discussion on the gap between answer-first prompt and logic-first prompt. To fill this gap, we designed a new benchmark method for testing language models based on this observation: Reasoning Order as Benchmark. This method reflects the consistency of the model’s reasoning logic by comparing the differences between the answer-first prompt and the logic-first prompt. Inspired from Chain-of-Thought (CoT) prompting (Wei et al. 2022) which enhances the performance of LLM reasoning by adding step-by-step guidance, we also designed Reflexive Prompting, a two-step prompting strategy. In the first step, we use both the answer-first prompt and the logic-first prompt to obtain two potentially different answers. In the second step, we analyze these two answers to derive the final answer. Experimental results demonstrate that this method improves the accuracy of reasoning tasks across different datasets and various LLMs. Furthermore, the accuracies achieved using different prompting strategies show strong correlation with the results of the Reasoning Order as Benchmark test result, proving the usability of this benchmarking approach.
@@
Related Works
Hallucination Problem
@@
Although the answers generated by LLMs may appear logically coherent, they are not always accurate and true. The hallucination problem causes these models to produce grammatically correct and contextually appropriate but false information (Bang et al. 2023; Agarwal et al. 2018), which may lead to disastrous consequences in ethically sensitive domains such as medicine, law consultation and autonomous driving.
Main causes of hallucinations focus on three aspects (Ye et al. 2023):
@@
  • Data Insufficient data (Wang et al. 2023) and noisy data (Yu et al. 2024) may limit the cross-modal feature alignment, thereby causing hallucination problems. Furthermore, (Yao et al. 2023) indicates that hallucination may be another view of adversarial examples.
@@
  • Knowledge Gap Knowledge gaps are usually ascribed to the variations in input formats between the pre-training and fine-tuning phases (Zheng, Huang, and Chang 2023). This knowledge gap creates complex challenges in balancing memory with retrieved evidence, which is seen as a passive defense mechanism to prevent the misuse of retrieval (Gao et al. 2022).
@@
  • Optimization Process Exposure bias between the training and testing stages have been demonstrated to lead to hallucinations within LLMs, particularly when generating lengthy responses (Wang and Sennrich 2020). In addition, sampling methods characterized by high uncertainty in the output phase, such as top-p and top-k, may also exacerbate the issue of hallucination (Lee et al. 2022).
Think Before You Speak
@@
The output stage of language models is typically a sequential recursive process: the $(K+1)^{th}$ output token is generated based on $K$ hidden vectors per layer, one vector preceding token. This mechanism is the most natural design choice when the Transformer was proposed by (Vaswani et al. 2017), ensuring the contextual coherence of the generated text, but it also means that the model does not consider potential special cases in the subsequent text when generating the preceding text.
@@
Research from (Goyal et al. 2023) investigated the possibility for the model to manipulate extra hidden vectors before outputting, for example, $K+10$ vectors during outputting $(K+1)^{th}$ token. By adding a sequence of dummy pause tokens into a decoder-only model’s input and ignore the model’s corresponding outputs until the last pause token is seen, the outputs are intentionally delayed to analyze the downstream task performances. Experiment results show that injecting delays in pretraining and fine-tuning during training stage outperforms the standard end-to-end training on wide variety of tasks, which also aligns with our hypothesis and results in this paper. Especially, pause-pretraining appears crucial for delays to help in downstream inference-time, indicating that a standard-pretrained model has strong biases that prevent it from fully understanding the benefits of inference-time delays.
Prompting Methods for Hallucination
@@
Due to the immense computational resources and data required to retrain large language models, this strategy is impractical for many researchers. Alternatively, modifying the prompting strategy to mitigate hallucination problems is more economically feasible. The following are some of the currently popular methods (Tonmoy et al. 2024):
@@
Prompting GPT-3 To Be Reliable (Si et al. 2022) proposed following prompting strategies for each aspect of reliability: (1)Prompting with randomly sampled examples from the source domain allows GPT-3 to generalize robustly on unseen domains and challenge examples; (2)Examples sampled from a balanced demographic distribution and natural language intervention reduce social biases; (3)Language model probabilities are calibrated to reflect accuracy; (4)Appending up-to-date knowledge can supplant GPT-3’s memorized knowledge or reasoning chains.
Chain-of-Verification (Dhuliawala et al. 2023) developed a CoT alike sequential approach namely Chain-of-Verification: for a given initial draft response, first plans verification questions to check its work, and then systematically answer those questions as a CoT procedure to gradually guide the model and finally produce an improved revised response.
Self-Reflecting Prompt (Ji et al. 2023) proposed an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation using a two-step loop: First acquire background knowledge and refine the knowledge, then answer the question and refine the answer referring to previous knowledge.
Self-Contradictory Hallucinations (Mündler et al. 2023) proposed a similar method to mitigate hallucination by a pair of contradictory prompt. The key difference is that the contradictory prompt is triggered initiatively by generating alternative sentence that aligns with original context within the same scope, while our method focuses more on passive contradictory caused by changing the answer-logic order, therefore no extra language extraction or inference models are needed to conduct such method.
Methods
@@
We noticed that in every response from the large language models with incorrect answer, the first sentence always presents the answer. Considering the sequential nature of language model outputs, at the point of generating the answer, the model has not yet seen the subsequent reasoning portion. Therefore, it may provide an incorrect answer based on similar causes identified in the training data, and then continue to generate reasoning to support this answer. To explore this phenomenon, we designed a set of new prompts for the question “Is 9.11 $>$ 9.9?”: (1) Prompts that instruct the model to first output the final result and then provide the reasoning, and (2) prompts that instruct the model to first output the reasoning and then the corresponding result. Surprisingly, we found that even if the LLMs can give out correct answer under original prompt, the issue of incorrect judgement may still occur when asked to generate an answer before reasoning logic, as in the comparison using Claude shown in Figure 2. Based on such observation, we propose the following hypotheses:
  • The sequential output of LLMs may lack of considerations for following outputs, sometimes cause hallucinations based on wrong pre-posing context;
@@
Reasoning Order as Benchmark
The result in Figure 2 confirms the correctness of our hypothesis that the sequential generation of language tokens by large language models, coupled with their inability to anticipate subsequent text during the generation of prior text, influences the accuracy of the model’s judgments, inducing a unique kind of illusionary phenomenon. From the perspective of natural language logic, the logical content and the final answer should remain consistent regardless of whether the reasoning logic is presented first or later. However, by specifying different output orders through prompt modification, the model’s outputs exhibit a form of inconsistency. This inconsistency indicates that at least one of the answers is experiencing a hallucination phenomenon, demonstrating the unreliability of the model’s reasoning. This insight has inspired us to propose a new benchmark method: Reasoning Order as Benchmark to measure the self-consistency of outputs from large language models.
The idea is quite simple: For each question, instead of its original prompt, generate a pair of “answer first” prompt and “logic first” prompt, and compare the consistency of the two results as benchmark. For example, to generate the “answer first” and “logic first” prompts, these two sentences are added at the end of the prompts: (1)Please give out the correct option in the first sentence and then give out the logic. (2)Please give out the reasoning logic first and then answer the question by selecting the options. The detailed benchmark method can be briefly illustrated in Algorithm 1.
Clearly, a robust language model should behave like a human, producing the same output for paraphrased queries that do not alter the underlying meaning. Consequently, evaluating the semantic stability of the model under such variations can indirectly validate its accuracy and robustness. This testing procedure is somewhat analogous to zero-shot adversarial attack-based evaluation methods (Shayegani et al. 2023), but it holds an advantage derived from Hypothesis 2: a reflexive prompt pair triggers adversarial attacks more efficiently. By sacrificing a portion of accuracy, this approach reduces testing time and cost, as it does not require additional computational resources to generate adversarial examples.
@@
In reflexive prompting, the LLM query process is transformed from a single-step direct inquiry into a two-step procedure. Similarly, in the first step, we generate the “answer first” and “logic first” prompt variants for the original query and get the corresponding results. In the second step, the original question and the two results are fed into the LLM altogether as a reflexive prompt for the final decision. The query process can be illustrated in Algorithm 2. An example of a reflexive prompt is shown below:
@@
Experiments
To evaluate the effectiveness of the reasoning order as benchmark as well as the reflexive prompting strategy, we designed the following experiments testing a variety of reasoning datasets on different LLMs. Experiment codes and results are available in https://github.com/XieZikai/ReflexivePrompting.
Claude Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
@@
Gemini Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
@@
Llama Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
@@
GPT Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
@@
Claude 0.870 0.897 0.884 0.902
Gemini 0.658 0.807 0.786 0.830
Llama 0.759 0.744 0.814 0.888
GPT 0.781 0.791 0.826 0.855
@@
Pearson Raw Prompt Answer First Logic First (0-shot CoT) Answer Ensemble Reflexive Prompt
@@
Preliminary Experiment
As noted above, one key hypothesis behind the idea of this project is that the answer-first and logic-first prompt pair can efficiently generate contradictory completions, so that the LLM may collect knowledge from both sides and infer the correct one. To verify this hypothesis as a sanity check, we compared our approach with the self-contradiction trigger algorithm proposed in (Mündler et al. 2023) in terms of the efficiency of triggering the contradiction. We follow the same prompt settings used in (Mündler et al. 2023), adding these two extra descriptions to form answer-first and logic-first strategies:
Please give me the logic/answer then give me the answer/logic.
We chose ChatGPT with GPT-3.5 turbo to carry out the experiment. With only 2 trials of prompt generating (answer-first and logic-first strategies), 7 out of 30 test questions triggered contradictions, achieving 23.3% self-contradiction trigger rate, which outperforms 17.7% reported from (Mündler et al. 2023). Although the experiment scale is rather limited and the experimental conditions are not entirely identical, considering that our approach is a noniterative, zero-shot method without additional knowledge, the result still validates the correctness of our hypothesis 2: Answer-first and logic-first prompt pairs may efficiently trigger the hallucination of self-inconsistency.
Experimental Settings
@@
  • Measureing Masssive Multitask Language Understanding(MMLU) (Hendrycks et al. 2020) is a multitask reasoning dataset, covering 57 tasks including elementary mathematics, US history, computer science, law and more. The validation set is chosen for the experiments, which contains 1,531 multiple choice questions.
  • TruthfulQA (Lin, Hilton, and Evans 2021) is a benchmark to measure whether a language model is truthful in generating answers that humans would answer falsely due to a false belief or misconception. The dataset comprises 817 questions that span 38 categories, including health, law, finance, and politics.
  • LogiQA (Liu et al. 2020) is an expert-written question dataset for testing human logical reasoning, which consists of 8,678 QA instances, covering multiple types of deductive reasoning. Due to a limited budget, only the first 1,000 questions are used in the experiments.
  • BigBench (bench authors 2023) contains 204 different tasks, covering areas such as linguistics, childhood development, math, common-sense reasoning and more. The multiple choice task “ruin_names” with 447 questions is chosen to evaluate the prompt strategies, which focus on selecting the humorous edit that ’ruins’ the input movie or musical artist name.
Models To conduct our experiments, we selected four commonly used large language models available on the market as test subjects, employing unadorned raw prompts, answer-first prompts, and logic-first prompts as baselines. These models are: GPT-4o-mini (Achiam et al. 2023), Llama-3.1-70b (Vavekanand and Sam 2024), Claude-3.5-sonnet (Amazon 2024) and Gemini-1.5-flash (Reid et al. 2024). In addition, it seems like an unfair comparison between reflexive prompt and the former prompt strategies for the extra LLM evaluations introducing extra model inferences and cost. Therefore, the ensemble model of three strategies voting is also added as a baseline.
Chain-of-thought prompting seems like a natural baseline for comparison. However, the logic-first prompting can be considered as equivalent to zero-shot CoT prompting (please see case study Figure 4 and Figure 4 for detail), and few-shot CoT prompting requires the design of specialized decomposition processes for each question, effectively introducing additional information. Therefore, we consider it unnecessary to compare it directly with few-shot CoT prompting.
Experimental Results

### v4

@@ Abstract
Order Matters in Hallucination: Reasoning Order as Benchmark and Reflexive Prompting for Large-Language-Models
Zikai Xie University of Science and Technology of China zikaix@ustc.edu.cn
1 Introduction
Large language models (LLMs) have revolutionized natural language processing (NLP), surpassing traditional models through their vast parameter counts and extensive training data. Their powerful generative and reasoning abilities have enabled wide adoption across domains such as education (Kasneci et al., 2023), healthcare (Sallam, 2023), and finance (Wu et al., 2023). Landmark models like GPT-3 Brown et al. (2020) and GPT-4 Achiam et al. (2023) have fundamentally reshaped language understanding and generation. However, the rapid deployment of AI has raised concerns, particularly in high-stakes, ethically sensitive areas. The opaque nature of machine learning models makes their decisions difficult to interpret, undermining accountability and risk control. As Morley et al. (2020) notes, their probabilistic outputs hinder the establishment of causal relationships, leaving room for error and misuse.
Even in systems that attempt to explain their reasoning, issues like hallucinations persist Yao et al. (2023), limiting broader adoption. Hallucinations occur when models generate factually incorrect yet plausible-sounding responses. A striking example is the widespread failure of LLMs to correctly compare numbers—many incorrectly assert that 9.11$>$9.9. As shown in Figure 1, most models respond that 9.11 is greater, often justifying the error with flawed or even logically sound reasoning that nonetheless leads to the wrong conclusion.
@@ 1 Introduction
Although recent studies have demonstrated the impact of prompting strategies on mitigating hallucination problems across various fields Si et al. (2022); Mündler et al. (2023); Ji et al. (2023); Dhuliawala et al. (2023), there is a lack of research and discussion on the gap between answer-first prompt and logic-first prompt. To fill this gap, we designed a new benchmark method for testing language models based on this observation: Reasoning Order as Benchmark. This method reflects the consistency of the model’s reasoning logic by comparing the differences between the answer-first prompt and the logic-first prompt. Inspired from Chain-of-Thought (CoT) prompting Wei et al. (2022) which enhances the performance of LLM reasoning by adding step-by-step guidance, we also designed Reflexive Prompting, a two-step prompting strategy. In the first step, we use both the answer-first prompt and the logic-first prompt to obtain two potentially different answers. In the second step, we analyze these two answers to derive the final answer. Experimental results demonstrate that this method improves the accuracy of reasoning tasks across different datasets and various LLMs. Furthermore, the accuracies achieved using different prompting strategies show strong correlation with the results of the Reasoning Order as Benchmark test result, proving the usability of this benchmarking approach.
@@
2 Related Works
2.1 Hallucination Problem
@@
Although the answers generated by LLMs may appear logically coherent, they are not always accurate and true. The hallucination problem causes these models to produce grammatically correct and contextually appropriate but false information Bang et al. (2023); Agarwal et al. (2018), which may lead to disastrous consequences in ethically sensitive domains such as medicine, law consultation and autonomous driving.
Main causes of hallucinations focus on three aspects Ye et al. (2023):
@@
  • Data Insufficient data Wang et al. (2023) and noisy data Yu et al. (2024) may limit the cross-modal feature alignment, thereby causing hallucination problems. Furthermore, Yao et al. (2023) indicates that hallucination may be another view of adversarial examples.
@@
  • Knowledge Gap Knowledge gaps are usually ascribed to the variations in input formats between the pre-training and fine-tuning phases Zheng et al. (2023). This knowledge gap creates complex challenges in balancing memory with retrieved evidence, which is seen as a passive defense mechanism to prevent the misuse of retrieval Gao et al. (2022).
@@
  • Optimization Process Exposure bias between the training and testing stages have been demonstrated to lead to hallucinations within LLMs, particularly when generating lengthy responses Wang and Sennrich (2020). In addition, sampling methods characterized by high uncertainty in the output phase, such as top-p and top-k, may also exacerbate the issue of hallucination Lee et al. (2022).
2.2 Think Before You Speak
@@
The output stage of language models is typically a sequential recursive process: the $(K+1)^{th}$ output token is generated based on $K$ hidden vectors per layer, one vector preceding token. This mechanism is the most natural design choice when the Transformer was proposed by Vaswani et al. (2017), ensuring the contextual coherence of the generated text, but it also means that the model does not consider potential special cases in the subsequent text when generating the preceding text.
@@
Research from Goyal et al. (2023) investigated the possibility for the model to manipulate extra hidden vectors before outputting, for example, $K+10$ vectors during outputting $(K+1)^{th}$ token. By adding a sequence of dummy pause tokens into a decoder-only model’s input and ignoring the model’s corresponding output until the last pause token is seen, the outputs are intentionally delayed to analyze the downstream task performances. The results of the experiment show that injecting delays in pre-training and fine-tuning during the training stage outperforms standard end-to-end training on a wide variety of tasks, which also aligns with our hypothesis and results in this paper. Especially, pause-pretraining appears crucial for delays to help in downstream inference time, indicating that a standard-pretrained model has strong biases that prevent it from fully understanding the benefits of inference-time delays.
2.3 Prompting Methods for Hallucination
@@
Due to the immense computational resources and data required to retrain large language models, this strategy is impractical for many researchers. Alternatively, modifying the prompting strategy to mitigate hallucination problems is more economically feasible. The following are some of the currently popular methods Tonmoy et al. (2024):
@@
Prompting GPT-3 To Be Reliable Si et al. (2022) proposed following prompting strategies for each aspect of reliability: (1)Prompting with randomly sampled examples from the source domain allows GPT-3 to generalize robustly on unseen domains and challenge examples; (2)Examples sampled from a balanced demographic distribution and natural language intervention reduce social biases; (3)Language model probabilities are calibrated to reflect accuracy; (4)Appending up-to-date knowledge can supplant GPT-3’s memorized knowledge or reasoning chains.
Chain-of-Verification Dhuliawala et al. (2023) developed a CoT alike sequential approach namely Chain-of-Verification: for a given initial draft response, first plans verification questions to check its work, and then systematically answer those questions as a CoT procedure to gradually guide the model and finally produce an improved revised response.
Self-Reflecting Prompt Ji et al. (2023) proposed an interactive self-reflection methodology that incorporates knowledge acquisition and answer generation using a two-step loop: First acquire background knowledge and refine the knowledge, then answer the question and refine the answer referring to previous knowledge.
Self-Contradictory Hallucinations Mündler et al. (2023) proposed a similar method to mitigate hallucination using a pair of contradictory prompts. The key difference is that the contradictory prompt is triggered initiatively by generating alternative sentence that aligns with original context within the same scope, while our method focuses more on passive contradictory caused by changing the answer-logic order, therefore no extra language extraction or inference models are needed to conduct such method.
3 Methods
@@
We noticed that in every response from large language models with incorrect answer, the first sentence always presents the answer. Considering the sequential nature of language model outputs, at the point of generating the answer, the model has not yet seen the subsequent reasoning portion. Therefore, it may provide an incorrect answer based on similar causes identified in the training data, and then continue to generate reasoning to support this answer. To explore this phenomenon, we designed a set of new prompts for the question “Is 9.11 $>$ 9.9?”: (1) Prompts that instruct the model to first output the final result and then provide the reasoning, and (2) prompts that instruct the model to first output the reasoning and then the corresponding result. Surprisingly, we found that even if the LLMs can give out correct answer under original prompt, the issue of incorrect judgement may still occur when asked to generate an answer before reasoning logic, as in the comparison using Claude shown in Figure 2. Based on such observation, we propose the following hypotheses:
  • LLMs may generate outputs sequentially without considering future responses, which can sometimes lead to hallucinations based on incorrect or misleading prior context;
@@
To test these hypotheses in a more systematic and scalable manner, we further designed an evaluation framework that quantifies the consistency and reliability of LLM outputs under different prompting orders. This framework, which we refer to as Reasoning Order as Benchmark, enables us to assess how the ordering of reasoning and answer affects the model’s behavior across diverse datasets and models.
3.1 Reasoning Order as Benchmark
The result in Figure 2 supports our hypothesis: the sequential generation of tokens by large language models, without the ability to foresee future content, affects the accuracy of their judgments. This limitation can induce a unique form of hallucination. In principle, the reasoning process and the final answer should remain consistent, regardless of whether the explanation appears before or after the answer. However, when we modify the prompt to alter the order of reasoning and answer, the model’s responses become inconsistent. This inconsistency suggests that at least one of the outputs is a hallucination, highlighting a weakness in the model’s reasoning capability. Based on this observation, we propose a new evaluation method, Reasoning Order as Benchmark, to assess the self-consistency of large language model outputs.
The idea is quite simple: For each question, instead of its original prompt, generate a pair of “answer first” prompt and “logic first” prompt and compare the consistency of the two results as a benchmark. For example, to generate the “answer first” and “logic first” prompts, these two sentences are added at the end of the prompts: (1)Please give out the correct option in the first sentence and then the logic. (2)Please give out the reasoning logic first and then answer the question by selecting the options. The detailed benchmark method can be briefly illustrated in Algorithm 1.
Clearly, a robust language model should behave like a human, producing the same output for paraphrased queries that do not alter the underlying meaning. Consequently, evaluating the semantic stability of the model under such variations can indirectly validate its accuracy and robustness. This testing procedure is somewhat analogous to zero-shot adversarial attack-based evaluation methods Shayegani et al. (2023), but it holds an advantage derived from Hypothesis 2: a reflexive prompt pair triggers adversarial attacks more efficiently. By sacrificing a portion of accuracy, this approach reduces testing time and cost, as it does not require additional computational resources to generate adversarial examples.
@@ 3.2 Reflexive Prompting
In reflexive prompting, the LLM query process is transformed from a single-step direct inquiry into a two-step procedure. Similarly, in the first step, we generate the “answer first” and “logic first” prompt variants for the original query and get the corresponding results. In the second step, the original question and the two results are fed into the LLM as a reflexive prompt for the final decision. The query process can be illustrated in Algorithm 2. An example of a reflexive prompt is shown below:
@@
4 Experiments
To evaluate the effectiveness of the reasoning order as benchmark as well as the reflexive prompting strategy, we designed the following experiments testing a variety of reasoning datasets on different LLMs.
4.1 Preliminary Experiment
As noted above, one key hypothesis behind the idea of this project is that the answer-first and logic-first prompt pair can efficiently generate contradictory completions, so that the LLM may collect knowledge from both sides and infer the correct one. To verify this hypothesis as a sanity check, we compared our approach with the self-contradiction trigger algorithm proposed in Mündler et al. (2023) in terms of the efficiency of triggering the contradiction. We follow the same prompt settings used in Mündler et al. (2023), adding these two extra descriptions to form answer-first and logic-first strategies:
Please give me the logic/answer then give me the answer/logic.
We chose ChatGPT with GPT-3.5 turbo to carry out the experiment. With only 2 trials of prompt generating (answer-first and logic-first strategies), 7 out of 30 test questions triggered contradictions, achieving 23.3% self-contradiction trigger rate, which outperforms 17.7% reported from Mündler et al. (2023). Although the experiment scale is rather limited and the experimental conditions are not entirely identical, considering that our approach is a noniterative, zero-shot method without additional knowledge, the result still validates the correctness of our hypothesis 2: Answer-first and logic-first prompt pairs may efficiently trigger the hallucination of self-inconsistency.
Claude-3.5-sonnet Raw Prompt Answer First Logic First Answer Ensemble Reflexive Prompt
@@ 4.1 Preliminary Experiment
Gemini-1.5-flash Raw Prompt Answer First Logic First Answer Ensemble Reflexive Prompt
@@
Llama-3.1-70b Raw Prompt Answer First Logic First Answer Ensemble Reflexive Prompt
@@
GPT-4o-mini Raw Prompt Answer First Logic First Answer Ensemble Reflexive Prompt
@@
Claude-3.5-sonnet 0.870 0.897 0.884 0.902
Gemini-1.5-flash 0.658 0.807 0.786 0.830
Llama-3.1-70b 0.759 0.744 0.814 0.888
GPT-4o-mini 0.781 0.791 0.826 0.855
@@
Pearson Raw Prompt Answer First Logic First Answer Ensemble Reflexive Prompt
@@
4.2 Experimental Settings
@@
  • Measureing Masssive Multitask Language Understanding(MMLU) Hendrycks et al. (2020) is a multitask reasoning dataset, covering 57 tasks including elementary mathematics, US history, computer science, law and more. The validation set is chosen for the experiments, which contains 1,531 multiple choice questions.
  • TruthfulQA Lin et al. (2021) is a benchmark to measure whether a language model is truthful in generating answers that humans would answer falsely due to a false belief or misconception. The dataset comprises 817 questions that span 38 categories, including health, law, finance, and politics.
  • LogiQA Liu et al. (2020) is an expert-written question dataset for testing human logical reasoning, which consists of 8,678 QA instances, covering multiple types of deductive reasoning. Due to a limited budget, only the first 1,000 questions are used in the experiments.
  • BigBench bench authors (2023) contains 204 different tasks, covering areas such as linguistics, childhood development, math, common-sense reasoning and more. The multiple choice task “ruin_names” with 447 questions is chosen to evaluate the prompt strategies, which focus on selecting the humorous edit that ’ruins’ the input movie or musical artist name.
Models To conduct our experiments, we selected four commonly used large language models available on the market as test subjects, employing unadorned raw prompts, answer-first prompts, and logic-first prompts as baselines. These models are: GPT-4o-mini Achiam et al. (2023), Llama-3.1-70b Vavekanand and Sam (2024), Claude-3.5-sonnet Amazon (2024) and Gemini-1.5-flash Reid et al. (2024). In addition, it seems like an unfair comparison between reflexive prompt and the former prompt strategies for the extra LLM evaluations introducing extra model inferences and cost. Therefore, the ensemble model of three strategies voting is also added as a baseline.
Chain-of-thought prompting seems like a natural baseline for comparison. However, the logic-first prompting can be considered as equivalent to zero-shot CoT prompting (please see case study Figure A.2 and Figure A.2 for detail), and few-shot CoT prompting requires the design of specialized decomposition processes for each question, effectively introducing additional information. Therefore, we consider it unnecessary to compare it directly with few-shot CoT prompting.
4.3 Experimental Results